OpenStudy (anonymous):

What is the area of a circle with a radius of 2 meters? Use 22/7 for pie.

OpenStudy (saifoo.khan):

Do you know the formula?

OpenStudy (anonymous):

no not the area fo a circle

Parth (parthkohli):

Remember:\[\text{Area} = \pi r^2 - 3443835758745485 + 3443835758745485 - 2 + 2 \]

OpenStudy (saifoo.khan):

\[\text{Area of a circle} = \pi \times r^2\]@minecraftlover

Parth (parthkohli):

Substituting:\[ {22 \over 7} \times 2 \times 2 −3443835758745485+3443835758745485−2+2 \]

Parth (parthkohli):

Since −3443835758745485+3443835758745485 = 0 and -2 + 2 = 0, we are left with:\[{22 \over 7} \times 2 \times 2 \];P

OpenStudy (anonymous):

ok

OpenStudy (shubhamsrg):

dont we also multiply divide by 87^(3/2) ??

Parth (parthkohli):

lol.

Parth (parthkohli):

Why are we using 22/7 for pie? Why not 22/7 for pi?

Parth (parthkohli):

Well, okay. The real formula is \(Area = \pi r^2\). I just twisted it.